OpenStudy (anonymous):

help me please if there are some persons and dogs who are total 36 head and 104 legs then how many legs dogs have if 4x+2y=104

OpenStudy (bibby):

That looks like the equation for legs where x is dogs and y is humans. 4 legs to a dog 2 to a human Can you post the original question so I'm sure I'm headed the right way?

OpenStudy (anonymous):

yes there are 2 of peoples and 4 are dogs

OpenStudy (bibby):

Use full sentences so that the meaning isn't lost. as I understand it there are 36 heads and 104 legs in total we have the equation for legs, and both dogs AND humans have ONE head that means in total there are 36 humans and dogs

OpenStudy (anonymous):

In the dog show yesterday, there were some dogs and men standing in the middle of the arena. Mina counted a total of 36 heads and 104 legs. How many dogs were standing in the arena?

OpenStudy (bibby):

36 heads means that x+y =36 now we can get this in terms of one or the other By that I mean: y=36-x OR x = 36-y Then we can go back to the original equation 4x+2y=104 and replace one of the variables.